OpenStudy (anonymous):

a man weighs 750N on the earth surface.what would be his weight when standing on the moon? the masses of the earth and the moon are respectively 5*10^24kg and 7.36*10^22kg. their radii are respectively6.37*10^3km and1.74*10^3km

OpenStudy (anonymous):

We have Weight on earth= G*(Mass of earth).m/(radius of earth)^2=750 Weight on earth=G*(Mass of moon)*m/(radius of moon)^2=y Where m is the mass of the man ang G is a universal constant for gravitation from the equation 1 we calculate the mass of man(m) & put the value in 2nd equation to find out the weight on moon y.